if a=4/3 cot^2 30 +3 sin^2 60 -3/4 tan^2 30 and b=3 tan^2 45 +cos 0 -cot 90 then logb(a) is

Grade:12

2 Answers

Chetan Mandayam Nayakar
13 years ago

a=(4/3)*3 + (3/4) - (3/4)*(1/3)=(9/2),b=3+1-0=4,logb(a)=log4(9/2)

shubhra pathak
13 years ago

a=(4/3)*3+(3/4)^2 -(3/4)*(1/3)   b=3+1-0

logb(a)= log4(69/4)